Solve the system of equations. (If the system is dependent, enter a general solution in terms of c.
{2x− y+ z = 12
2y− 3z=−16
3y+ 2z= 2
(x, y, z) =

Answer by Alan3354(69443)


2x - y + z = 12
    2y - 3z = -16
    3y + 2z = 2
------
    6y - 9z = -48  ---- 2nd eqn times 3
    6y + 4z = 4    ---- 3rd eqn times 2
----------------------------- Subtract
-13z = -52
z = 4
==============
y = -2
=======
x = 5